OpenStudy (anonymous):

The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 480 feet. Show all steps in your work.

ganeshie8 (ganeshie8):

nice problem :) you're going to find the acceleration due to gravity !! good luck !

ganeshie8 (ganeshie8):

\(\large d = 0.5gt^2 \)

ganeshie8 (ganeshie8):

to solve, is to isolate so just isolate \(g\) by dividing \(0.5t^2\) both sides

OpenStudy (anonymous):

Sooooooo? will it be um , 0.5t2/g

ganeshie8 (ganeshie8):

\(\large d = 0.5gt^2\) divide 0.5t^2 both sides \(\large \frac{d}{0.5t^2} = g\) \(\large g = \frac{d}{0.5t^2}\) -------------------(1)

ganeshie8 (ganeshie8):

thats ur equation for finding g

OpenStudy (anonymous):

You're sooo sexy. ^_^

OpenStudy (anonymous):

um? @TheTitan

OpenStudy (anonymous):

2.The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (anonymous):

someone please help me